A circle contains two tangent lines with equations X=5 and Y=7.What would be the coordinates of the centre?

Grade:12

1 Answers

PRIYANKA TATA ASKIITIAN
13 years ago

since we are given two tangents to the circle , then the centre will lie on their angular bisectors of the two tangents.

the angular bisectors will be x-5=y-7 and x-5=-(y-7).

=> y=x+2 and y=12-x

then the cordinates of the centre can be (x,x+2) and (x,12-x)... putting the values from given options,you can get the centre value.

plz approve my answer..!

priyanka tata